# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 7. Asymptotics of the $\mathcal{K}$-reduced partition function on log Fano curves
In this section we prove Theorem 1.7. To this end it will be convenient to replace the basis of $H^{0}(X,-kK_{(X,\Delta)}),$ appearing in the definition of $\nu^{(N)},$ by the basis of monomials $1,z,z^{2},\dots$ induced by standard identification of $kK_{(X,\Delta)}$ with $kV\mathcal{O}(1)\rightarrow\mathbb{P}^{1}.$ This amounts to adding an appropriate constant to the corresponding Mabuchi functional $\mathcal{M}$ of $(X,\Delta)$ to get a new functional that we denote by $\mathcal{M}_{\mathbb{Z}}$ (the subscript indicates the arithmetic interpretation of this setup, discussed in Section 8). The inf of $\mathcal{M}_{\mathbb{Z}}$ was explicitly computed in [3, Thm 6.1], when $\Delta$ is any divisor on $\mathbb{P}^{1}$ with support $\{0,1,\infty\},$ assuming that $(X,\Delta)$ is K-stable (in particular, $\mathcal{G}$ is then trivial). In a nutshell, we will reduce the proof of Theorem 1.7, where $\mathcal{G}$ is non-trivial to the computation in [3, Thm 6.1], by showing that breaking the $\mathcal{G}$-symmetry corresponds, effectively, to perturbing the given K-polystable log Fano curve $(X,\Delta)$ to a K-stable log Fano curve $(X,\Delta_{N})$.
@@
On any log Fano curve $(X,\Delta)$ with vanishing Futaki character, $g^{*}\nu_{N}=\nu_{N}$ for any $g\in\mathcal{G}.$
@@
We can express $(X,\Delta)=(\mathbb{P}^{1},\Delta_{w})$ for $w\in[0,1[$ (see the discussion proceeding Theorem 2.11). Fix a $\mathbb{C}^{*}$–action preserving $\Delta_{w}.$ Then the corresponding moment polytope $P_{w}$ is symmetric (P=[1+w,1w])(P=[-1+w,1-w]) and thus $\boldsymbol{b}_{k}=0$ in Prop 6.20, which implies the lemma. ∎
@@
By Prop 6.14, $Z_{N}^{\mathrm{c}}$ (defined by formula 6.4) may be expressed as
@@
Note that $\eta$ defines a $\mathcal{G}$–invariant measure on $X^{3}$ (this also follows from Lemma 7.1, applied to $N=3).$ Thus formula 7.3 recovers the $\mathcal{G}$–invariance of $\nu^{(N)}$ in Lemma 6.9. Moreover, it follows from the $\mathcal{G}$–invariance of $\eta$ that the contracted form $\eta^{\mathrm{c}}$ is a constant. In fact, taking $e_{1},e_{2}$ and $e_{3}$ as the standard generators of the Lie algebra $\mathfrak{sl}(2,\mathbb{C})$ corresponding to the holomorphic vector fields $\partial/\partial z,$ $z\partial/\partial z$ and $z^{2}\partial/\partial z$ on $\mathbb{C}$ a direct computation gives
@@
Moreover, using that $y_{i}$ is $\mathcal{G}$–invariant we have $\delta_{e_{i}}dy_{j}=\delta_{Je_{i}}dy_{j}=0,$ in the notation of Section 6.4.1. Thus, accepting formula 7.3, we conclude that
@@
which implies the desired formula 7.2 (with $C=1),$ since $y_{i}$ is $\mathcal{G}$–invariant.
Finally, as next described, the deferred proof of formula 7.3 follows from a direct, but somewhat unwieldy calculations (the argument is guided by the simpler case when $\Delta$ is nontrivial and $\mathcal{G}=\mathbb{C}^{*},$ described in Section 7.2). First note that since $V=2$ and $N=2k+1$ [18, Lemma 4.3] yields
@@
which concludes the proof of formula 7.3.
7.1.2. Estimating the distortion $f_{N}$ between $Z_{\boldsymbol{m}_{N}}$ and $Z_{N}^{\mathrm{c}}$
We next show that there exists a positive constant $C$ such that function $f_{N}$ in formula 6.15 satisfies $N^{-3}C\leq f_{N}\leq C$ when $\boldsymbol{m}_{N}=0.$ By symmetry, it will be enough to show that
@@ 7.1.2. Estimating the distortion $f_{N}$ between $Z_{\boldsymbol{m}_{N}}$ and $Z_{N}^{\mathrm{c}}$
Denote by $P$ the convex polytope in $\mathbb{R}^{N}$ defined by the constraints on $p_{i}$ above. Since the function to be maximized is convex its maximum is attained at the extremal points $(p_{1},\dots,p_{N})$ of $P.$ We claim that these satisfy $p_{i}\in\{0,\pm 1\}.$ It then readily follows that the maximum in question is attained when $p_{i}=0$ for one index $i$ and the remaining N1N-1 indices are split into two groups of (N1)/2(N-1)/2 members satisfying $p_{i}=-1$ and $p_{i}=1,$ respectively, which proves the lower bound in 7.4. Finally, to prove the previous claim, consider the hypersimplex $\Delta_{N/2}$ obtained by replacing $p_{i}$ with $t_{i}\coloneqq(p_{i}+1)/2$:
@@ 7.1.3. Computing the large-$N$ limit of $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
As explained in [3], the general integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ defined in formula 7.1, can be viewed as an arithmetic partition function on $(\mathbb{P}^{1})^{N}$ corresponding to the log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}}),$ where $\Delta^{\boldsymbol{w}}$ denotes the divisor supported on the three points $0,1$ and $\infty$ with weights $w_{1},w_{2}$ and $w_{3},$ respectively (in this case the corresponding automorphism group $\mathcal{G}$ is trivial). By [3, Thm 5.3], given any log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ that is K-stable (or, equivalently, Gibbs stable), the corresponding partition function on $(\mathbb{P}^{1})^{N}$ satisfies
@@
where $\mathcal{M}_{\mathbb{Z}}$ denote the arithmetic Mabuchi functional attached to $(\mathbb{P}^{1},0).$ As pointed out in [3, Section 6.2], it follows from [53, Formula B.9] (and [60, formula 3.1] and [92, Section 2.3]) that for any weight vector $\boldsymbol{w}\in]0,1[^{3}$ such that $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ is a log Fano curve that is K-stable, i.e. such that
@@
(as follows from formula 7.8 applied to $V=0$ and any $N\geq 2).$ When $\boldsymbol{w}$ is independent of $N$ it follows from elementary arguments, as observed in [3], that $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}}$ converges, as $N\rightarrow\infty,$ towards
@@
Combining this convergence with the convergence 7.5 gives the following explicit formula
@@
which is shown in [3] to be valid also when $(X,\Delta)$ is merely K-polystable, using a continuity argument with respect to $\boldsymbol{w}.$ In particular, the formula applies to $\Delta=0.$ Now apply the explicit formula 7.10 to the case of the weight vector $\boldsymbol{w}_{N}$ in formula 7.6. Note that
@@
where the error terms are positive. Hence, Stirling’s approximation and the fact that $\Gamma(x)$ has a simple pole with residue 1 at $x=0$ leads to the following logarithmic asymptotics of the first factor in formula 7.8:
@@
using that the Gamma function has a simple pole at $x=0$ and that $x\log x-x$ is a primitive of $\log x.$ This proves the asymptotics 7.11. The nominator in the second factor in formula 7.8 is handled in essentially the same manner, giving
@@
for $w_{i}=0$ and $V=2.$ The proof of formula 7.6 is thus concluded by invoking formula 7.10.
@@
To prove this formula recall that, by [18, Lemma 4.3], we may express
@@
for some real number $r$ and a direct computation reveals that $r=0.$ Alternatively, using that both $\nu^{(N)}$ and the second factor in the right hand side of the previous equation is $\mathcal{G}$–invariant (i.e. $\mathbb{C}^{*}$–invariant) forces $r=0.$ This proves formula 7.13.
@@ 8. Relations to arithmetic geometry
(8.1) $\lim_{N\rightarrow\infty}N^{-1}\log\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}=\inf\mathcal{M}_{\mathbb{Z}}.$
In fact, this conjecture is equivalent to the convergence in Conjecture 1.6, as follows readily from the arithmetic Hilbert–Samuel theorem (by the argument in the proof of [3, Thm 5.3]). One drawback of Conjecture 1.6 – from the arithmetic point of view – is that left hand side in formula 8.1 depends on non-arithmetic data (the choice of the group $\mathcal{K}$ and the metric $\phi_{0}$).). This can be remedied if the quantized Futaki character vanishes for $k$ sufficiently large (Section 6.5) and the Lie algebra $\mathfrak{g}$ of the $\mathcal{G}$ is the complexification of a natural $\mathbb{Z}$-module $\mathfrak{g_{\mathbb{Z}}}$ attached to $(\mathcal{X},\mathcal{D})$:
@@
realized as periods. The following conjecture generalizes [18, Conjecture 5.2] (see also [3, Section 5.2]) to the case when $\mathcal{G}$ is nontrivial:
@@
By Prop 6.14, the two previous conjectures are equivalent if the $f_{N}$ defined in formula 6.15 satisfies $\log f_{N}=o(N^{-1}).$ We expect this to be the case, in general (the case when $X=\mathbb{P}^{1}$ was established in the proof of Theorem 1.7).
@@ 8.1. Examples
The condition 8.2 is, for example, satisfied when $\mathcal{X}$ is smooth over $\mathbb{Z},$ if $\mathfrak{g_{\mathbb{Z}}}$ is taken as the relative tangent bundle $T_{\mathcal{X}/\mathbb{Z}}$ of $\mathcal{X}$ over $\mathbb{Z}.$ This is the case if $\mathcal{X}$ is homogeneous over $\mathbb{Z},$ i.e. $\mathcal{X}=G/P,$ where $G$ is a semisimple Chevalley group over $\mathbb{Z}$ and $P$ a parabolic subgroup of $G$ [66, Section 2], or if $\mathcal{X}$ is the canonical model over $\mathbb{Z}$ of a toric Fano manifold $X$ [81, Section 2]. By Prop 6.20, the assumptions of Conjecture 8.2 are then satisfied (for all $k)$ if, in the toric case, we also assume that the quantized barycenters vanish. It should be stressed that in these cases, the arithmetic data that go into the definition of the periods $Z_{N}(\mathcal{X},\mathcal{D})$ is, essentially, combinatorial. One is thus led to ask if $Z_{N}(\mathcal{X},\mathcal{D})$ can be computed explicitly in these cases. For example, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ and $\mathcal{D}$ is trivial, combining formula 7.1 and formula 7.8 shows that $Z_{N}(\mathcal{X},\mathcal{D})$ is a product of Gamma-factors at rational numbers (up to factors of $\pi)$ with rational exponents, when $N\geq 5$:
@@
(and a similar formula holds when $N=4,$ by formula 7.9). Likewise, when $\mathcal{D}$ is the Zariski closure $\mathcal{D}_{w}$ in $\mathbb{P}_{\mathbb{Z}}^{1}$ of the divisor $\Delta$ over $\mathbb{Q}$ supported at $\{0,\infty\}$ with equal weights $w\in]0,\infty[,$ combining formula 7.12 and formula 7.8 gives, when $N\geq 3,$

### v3

@@ 8. Asymptotics of the $\mathcal{K}$-reduced partition function on log Fano curves
In this section we prove Theorem 1.9. To this end it will be convenient to replace the basis of $H^{0}(X,-kK_{(X,\Delta)}),$ appearing in the definition of $\nu^{(N)},$ by the basis of monomials $1,z,z^{2},\dots$ induced by standard identification of $kK_{(X,\Delta)}$ with $kV\mathcal{O}(1)\rightarrow\mathbb{P}^{1}.$ This amounts to adding an appropriate constant to the corresponding Mabuchi functional $\mathcal{M}$ of $(X,\Delta)$ to get a new functional that we denote by $\mathcal{M}_{\mathbb{Z}}$ (the subscript indicates the arithmetic interpretation of this setup, discussed in Section 9). The inf of $\mathcal{M}_{\mathbb{Z}}$ was explicitly computed in [4, Thm 6.1], when $\Delta$ is any divisor on $\mathbb{P}^{1}$ with support $\{0,1,\infty\},$ assuming that $(X,\Delta)$ is K-stable (in particular, $\mathcal{G}$ is then trivial). In a nutshell, we will reduce the proof of Theorem 1.9, where $\mathcal{G}$ is non-trivial to the computation in [4, Thm 6.1], by showing that breaking the $\mathcal{G}$-symmetry corresponds, effectively, to perturbing the given K-polystable log Fano curve $(X,\Delta)$ to a K-stable log Fano curve $(X,\Delta_{N})$.
@@
On any log Fano curve $(X,\Delta)$ with vanishing Futaki character, $g^{*}\nu_{N}=\nu_{N}$ for any $g\in\mathcal{G}.$ Moreover, for any log Fano curve 0 is a regular value of the moment map $\boldsymbol{m}_{N}$ if $N$ is odd and $N\geq 3.$
@@
We can express $(X,\Delta)=(\mathbb{P}^{1},\Delta_{w})$ for $w\in[0,1[$ (see the discussion proceeding Theorem 2.11). Fix a $\mathbb{C}^{*}$–action preserving $\Delta_{w}.$ Then the corresponding moment polytope $P_{w}$ is symmetric (P=[1+w,1w])(P=[-1+w,1-w]) and thus $\boldsymbol{b}_{k}=0$ in Prop 6.22, which implies the first part of the lemma. Alternatively, one can show by direct computation that $g^{*}\nu_{N}=\nu_{N}$ for any $g\in\mathcal{G}.$ To prove the last statement, first note that for any $N\geq 2,$ $\left\{\boldsymbol{m}_{N}=0\right\}$ is non-empty, since it intersects the $\mathcal{G}$-orbit of any configuration $(x_{1},...,x_{N})$ of mutually distinct points. Indeed, by [103, Section 3], a configuration $(x_{1},...,x_{N})$ is stable iff the inequality 2.10 is strict and by the general Kempf-Ness isomorphism [83], the $\mathcal{G}$-orbit of any polystable point intersects the zero-locus of the moment map (and stability implies polystability, in general). Finally, assume that $N$ is odd and $N\geq 3.$ Then $(x_{1},...,x_{N})$ is semi-stable iff it is stable. Indeed, when $N$ is odd the semi-stability inequality 2.10 holds iff it is a strict inequality. Hence, the semi-stable locus coincides with the stable locus, which is equivalent to 0 beeing a regular value of the moment map [83, Cor 3.10]. ∎
@@
By Prop 6.16, $Z_{N}^{\mathrm{c}}$ (defined by formula 6.6) may be expressed as
@@
Note that $\eta$ defines a $\mathcal{G}$–invariant measure on $X^{3}$ (this also follows from Lemma 8.1, applied to $N=3).$ Thus formula 8.3 recovers the $\mathcal{G}$–invariance of $\nu^{(N)}$ in Lemma 6.1. Moreover, it follows from the $\mathcal{G}$–invariance of $\eta$ that the contracted form $\eta^{\mathrm{c}}$ is a constant. In fact, taking $e_{1},e_{2}$ and $e_{3}$ as the standard generators of the Lie algebra $\mathfrak{sl}(2,\mathbb{C})$ corresponding to the holomorphic vector fields $\partial/\partial z,$ $z\partial/\partial z$ and $z^{2}\partial/\partial z$ on $\mathbb{C}$ a direct computation gives
@@
Moreover, using that $y_{i}$ is $\mathcal{G}$–invariant we have $\delta_{e_{i}}dy_{j}=\delta_{Je_{i}}dy_{j}=0,$ in the notation of Section 6.5.1. Thus, accepting formula 8.3, we conclude that
@@
which implies the desired formula 8.2 (with $C=1),$ since $y_{i}$ is $\mathcal{G}$–invariant.
Finally, as next described, the deferred proof of formula 8.3 follows from a direct, but somewhat unwieldy calculations (the argument is guided by the simpler case when $\Delta$ is nontrivial and $\mathcal{G}=\mathbb{C}^{*},$ described in Section 8.2). First note that since $V=2$ and $N=2k+1$ [19, Lemma 4.3] yields
@@
which concludes the proof of formula 8.3.
8.1.2. Estimating the distortion $f_{N}$ between $Z_{N,0}$ and $Z_{N}^{\mathrm{c}}$
We next show that there exists a positive constant $C$ such that function $f_{N}$ in formula 6.17 satisfies $N^{-3}C\leq f_{N}\leq C$ when $\boldsymbol{m}_{N}=0.$ By symmetry, it will be enough to show that
@@ 8.1.2. Estimating the distortion $f_{N}$ between $Z_{N,0}$ and $Z_{N}^{\mathrm{c}}$
Denote by $P$ the convex polytope in $\mathbb{R}^{N}$ defined by the constraints on $p_{i}$ above. Since the function to be maximized is convex its maximum is attained at the extremal points $(p_{1},\dots,p_{N})$ of $P.$ We claim that these satisfy $p_{i}\in\{0,\pm 1\}.$ It then readily follows that the maximum in question is attained when $p_{i}=0$ for one index $i$ and the remaining N1N-1 indices are split into two groups of (N1)/2(N-1)/2 members satisfying $p_{i}=-1$ and $p_{i}=1,$ respectively, which proves the lower bound in 8.4. Finally, to prove the previous claim, consider the hypersimplex $\Delta_{N/2}$ obtained by replacing $p_{i}$ with $t_{i}\coloneqq(p_{i}+1)/2$:
@@ 8.1.3. Computing the large-$N$ limit of $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
As explained in [4], the general integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ defined in formula 8.1, can be viewed as an arithmetic partition function on $(\mathbb{P}^{1})^{N}$ corresponding to the log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}}),$ where $\Delta^{\boldsymbol{w}}$ denotes the divisor supported on the three points $0,1$ and $\infty$ with weights $w_{1},w_{2}$ and $w_{3},$ respectively (in this case the corresponding automorphism group $\mathcal{G}$ is trivial). By [4, Thm 5.3], given any log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ that is K-stable (or, equivalently, Gibbs stable), the corresponding partition function on $(\mathbb{P}^{1})^{N}$ satisfies
@@
where $\mathcal{M}_{\mathbb{Z}}$ denote the arithmetic Mabuchi functional attached to $(\mathbb{P}^{1},0).$ As pointed out in [4, Section 6.2], it follows from [69, Formula B.9] (and [76, formula 3.1] and [110, Section 2.3]) that for any weight vector $\boldsymbol{w}\in]0,1[^{3}$ such that $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ is a log Fano curve that is K-stable, i.e. such that
@@
(as follows from formula 8.8 applied to $V=0$ and any $N\geq 2).$ When $\boldsymbol{w}$ is independent of $N$ it follows from elementary arguments, as observed in [4], that $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}}$ converges, as $N\rightarrow\infty,$ towards
@@
Combining this convergence with the convergence 8.5 gives the following explicit formula
@@
which is shown in [4] to be valid also when $(X,\Delta)$ is merely K-polystable, using a continuity argument with respect to $\boldsymbol{w}.$ In particular, the formula applies to $\Delta=0.$ Now apply the explicit formula 8.10 to the case of the weight vector $\boldsymbol{w}_{N}$ in formula 8.6. Note that
@@
where the error terms are positive. Hence, Stirling’s approximation and the fact that $\Gamma(x)$ has a simple pole with residue 1 at $x=0$ leads to the following logarithmic asymptotics of the first factor in formula 8.8:
@@
using that the Gamma function has a simple pole at $x=0$ and that $x\log x-x$ is a primitive of $\log x.$ This proves the asymptotics 8.11. The nominator in the second factor in formula 8.8 is handled in essentially the same manner, giving
@@
for $w_{i}=0$ and $V=2.$ The proof of formula 8.6 is thus concluded by invoking formula 8.10.
@@
To prove this formula recall that, by [19, Lemma 4.3], we may express
@@
for some real number $r$ and a direct computation reveals that $r=0.$ Alternatively, using that both $\nu^{(N)}$ and the second factor in the right hand side of the previous equation is $\mathcal{G}$–invariant (i.e. $\mathbb{C}^{*}$–invariant) forces $r=0.$ This proves formula 8.13.
@@ 9. Relations to arithmetic geometry
(9.1) $\lim_{N\rightarrow\infty}N^{-1}\log\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{0}^{(N)}=\inf\mathcal{M}_{\mathbb{Z}}.$
In fact, this conjecture is equivalent to the convergence in Conjecture 1.8, as follows readily from the arithmetic Hilbert–Samuel theorem (by the argument in the proof of [4, Thm 5.3]). One drawback of Conjecture 1.8 – from the arithmetic point of view – is that left hand side in formula 9.1 depends on non-arithmetic data (the choice of the group $\mathcal{K}$ and the metric $\phi_{0}$).). This can be remedied if the quantized Futaki character vanishes for $k$ sufficiently large (Section 6.6) and the Lie algebra $\mathfrak{g}$ of the $\mathcal{G}$ is the complexification of a natural $\mathbb{Z}$-module $\mathfrak{g_{\mathbb{Z}}}$ attached to $(\mathcal{X},\mathcal{D})$:
@@
realized as periods. The following conjecture generalizes [19, Conjecture 5.2] (see also [4, Section 5.2]) to the case when $\mathcal{G}$ is nontrivial:
@@
By Prop 6.16, the two previous conjectures are equivalent if the $f_{N}$ defined in formula 6.17 satisfies $\log f_{N}=o(N^{-1}).$ We expect this to be the case, in general (the case when $X=\mathbb{P}^{1}$ was established in the proof of Theorem 1.9).
@@ 9.1. Examples
The condition 9.2 is, for example, satisfied when $\mathcal{X}$ is smooth over $\mathbb{Z},$ if $\mathfrak{g_{\mathbb{Z}}}$ is taken as the relative tangent bundle $T_{\mathcal{X}/\mathbb{Z}}$ of $\mathcal{X}$ over $\mathbb{Z}.$ This is the case if $\mathcal{X}$ is homogeneous over $\mathbb{Z},$ i.e. $\mathcal{X}=G/P,$ where $G$ is a semisimple Chevalley group over $\mathbb{Z}$ and $P$ a parabolic subgroup of $G$ [86, Section 2], or if $\mathcal{X}$ is the canonical model over $\mathbb{Z}$ of a toric Fano manifold $X$ [99, Section 2]. By Prop 6.22, the assumptions of Conjecture 9.2 are then satisfied (for all $k)$ if, in the toric case, we also assume that the quantized barycenters vanish. It should be stressed that in these cases, the arithmetic data that go into the definition of the periods $Z_{N}(\mathcal{X},\mathcal{D})$ is, essentially, combinatorial. One is thus led to ask if $Z_{N}(\mathcal{X},\mathcal{D})$ can be computed explicitly in these cases. For example, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ and $\mathcal{D}$ is trivial, combining formula 8.1 and formula 8.8 shows that $Z_{N}(\mathcal{X},\mathcal{D})$ is a product of Gamma-factors at rational numbers (up to factors of $\pi)$ with rational exponents, when $N\geq 5$:
@@
(and a similar formula holds when $N=4,$ by formula 8.9). Likewise, when $\mathcal{D}$ is the Zariski closure $\mathcal{D}_{w}$ in $\mathbb{P}_{\mathbb{Z}}^{1}$ of the divisor $\Delta$ over $\mathbb{Q}$ supported at $\{0,\infty\}$ with equal weights $w\in]0,\infty[,$ combining formula 8.12 and formula 8.8 gives, when $N\geq 3,$